A balloon is filled with hydrogen at 26oC. The balloon expands until the pressure is equal to the barometric pressure of 750 Torr. The balloon rises to an altitude of 6000 ft, where the pressure is 600 Torr and the temperature is 14oC. What is the change in the volume of the balloon as it ascends to 6000 ft ?
1
Expert's answer
2013-03-13T08:24:43-0400
This question can be solved by using the ideal gas law:

PV = nRT

But even if you have two different conditions the amount is the same (R is const), so:

P1V1 = nRT1
n = P1V1/RT1
n = P2V2/RT2
P1V1/RT1 = P2V2/RT2
P1V1/T1 = P2V2/T2

If starting volume is V1:

T1 = 26 C
T2 = 14 C

P1 = 750 Torr
P2 = 600 Torr

750 Torr V1/26 C = 600 Torr V2/14 C
28.85 V1 = 42.86 V2

0,67*V1 = V2